OpenStudy (anonymous):

Find the point on the parabola x+y^2=0 that is closest to the point (o,-3)

OpenStudy (anonymous):

So just by looking at it you can tell that the vertex is \((0, 0)\).

OpenStudy (anonymous):

We also know it opens up to the left or right, because it is horizontal.

OpenStudy (anonymous):

Getting it in standard parabola form \(y^2 = 4px\) we get: \[ y^2 = 4\left(-\frac{1}{4}\right)x \]This means the focus is at \((-1/4, 0)\) and the directrix is \(x=1/4\)

OpenStudy (anonymous):